# FractalForensics: Proactive Deepfake Detection and Localization via Fractal Watermarks

### v1

@@ 3.4. Loss Functions
Table 2. Watermarked recovery rate against common image operations on CelebA-HQ.
@@
Additionally, we train the decoder for watermark robustness following
@@ 4.1. Implementation Details
We used the popular facial image datasets, CelebA-HQ (Karras et al., 2018) and LFW (Huang et al., 2012), for experiments. Specifically, we followed the official split of CelebA-HQ with 30,000 facial images for training, validation, and testing, and we used LFW with 5,749 facial identities for evaluating the generalizability. We trained two versions of FractalForensics for the 256 and 512 resolutions, with $N_{\textrm{img}}$ being 5 and 6, $N_{\textrm{wtm}}$ being 5 and 4, $N_{\textrm{rec}}$ being 5 and 4, and $N_{\textrm{dec}}$ being 4 and 3, respectively. For the loss weights, we set $\lambda_{\textrm{MSE}}=1$, $\lambda_{\textrm{LPIPS}}=0.5$, $\lambda_{\textrm{adv}}=0.01$, and $\lambda_{\textrm{dec}}=12$. During training, we adopted Identity, Jpeg, and Gaussian Noise as common image processing operations to boost watermark robustness. Deepfake manipulations are completely omitted during training under the black-box setting. In the testing phase, the semi-fragile watermarking algorithms are evaluated against common image processing operations including Identity, Jpeg, Gaussian Noise, Gaussian Blur, Median Blur, and Resize and Deepfake manipulations including SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), UniFace (Xu et al., 2022), E4S (Liu et al., 2023), and DiffSwap (Zhao et al., 2023b) for face swapping and StarGAN (Choi et al., 2020), StyleMask (Bounareli et al., 2023b), and HyperReenact (Bounareli et al., 2023a) for face reenactment. We trained the main watermarking framework at a learning rate of $0.002$ and trained the discriminator at $0.0004$. All experiments are conducted on 8 Tesla A100 GPUs.
@@ 4.2. Watermarking Performance on CelebA-HQ
In this section, we evaluated the watermarking performance on CelebA-HQ (Karras et al., 2018), with respect to the visual quality of the watermarked images and the watermark robustness and fragility with respect to common image operations and Deepfake manipulations.
Visual Quality. We conducted quantitative evaluations of the visual quality after embedding watermarks. Specifically, we compared our proposed FractalForensics to state-of-the-art semi-fragile watermarks for proactive defense against Deepfake and AIGC. As Table 1 reports, our method achieves the second-best performance with the most difficult bit-per-pixel rate, while the best WaterLo (Beuve et al., 2023) has watermarks with only 1 bit. In general, while all methods demonstrate reasonable LIPIS values, SepMark [^3] SepMark consists of a robust and a semi-fragile watermark, where we adopted the semi-fragile watermark for comparing robustness and fragility in this study. (Wu et al., 2023) and FaceSigns (Neekhara et al., 2024) both have PSNR values below 35, and EditGuard (Zhang et al., 2024) derives an unsatisfactory SSIM value below 0.85. As a result, our approach consistently maintains promising visual qualities at both resolutions for all evaluation metrics on visual quality.
Table 3. Watermarked recovery rate against Deepfake image manipulations on CelebA-HQ.
@@
Watermark Recovery Rate. We evaluated the watermark recovery rate of our approach against common image processing operations and Deepfake manipulations, compared to contrastive semi-fragile watermarking frameworks. As demonstrated in Table 2, while all algorithms stay robust as expected when applied with no operation (Identity), they each suffer drastic damping in watermark recovery rates when facing some image processing operations. In general, Gaussian Noise is the one that has brought the most trouble, causing the watermark recovery rates all below 65% for the comparative algorithms. Meanwhile, the two models that have demonstrated relatively better visual quality in Table 1, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024), also fail to maintain robustness regarding Jpeg, Gaussian Blur, Median Blur, and Resize operations. For our proposed FractalForensics, since each watermark matrix entry is represented by a 4-bit binary value, we computed the patch-wise watermark recovery rate as our official performance and used that for Deepfake detection and localization in later sections, which is harder than the bit-wise metric in preserving robustness. As a result, our method achieves the best performance on average with 99.55% and 99.73% at the 256 and 512 resolutions. In the last two columns, we also presented the bit-wise watermark recovery rates of our proposed FractalForensics as references, where reasonably higher values than the patch-wise ones can be observed as expected since the patch-wise watermark recovery rate is computed following a more strict rule.
@@
When tested against Deepfake manipulations, the watermark fragility is expected to be observed. In other words, there is expected to be an obvious gap between the watermark recovery rates for robustness and fragility, such that real and fake samples are distinguishable accordingly. For this experiment, we adopted five face swapping (Chen et al., 2020; Gao et al., 2021; Xu et al., 2022; Liu et al., 2023; Zhao et al., 2023b) and three face reenactment (Choi et al., 2020; Bounareli et al., 2023b, a) manipulations for evaluation. It can be observed from Table 3 that, the fragility has been favorably accomplished by WaterLo (Beuve et al., 2023), FaceSigns (Neekhara et al., 2024), EditGuard (Zhang et al., 2024), and our proposed FractalForensics, while SepMark (Wu et al., 2023) demonstrates robust watermark recovery ability from images generated by SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), and E4S (Liu et al., 2023). Meanwhile, FractalForensics is observed to demonstrate promising fragility at both resolutions with average watermark recovery rates of 50.05% and 39.27%.
@@
With a closer look at Table 2 and Table 3, since there exist watermark recovery rates against common image operations that are lower than those against Deepfake manipulations for the comparative models (e.g., Sepmark has 99.94% for InfoSwap but 54.47% for Gaussian Noise), it becomes impossible to distinguish real and fake based on robustness and fragility. As a result, although our FractalForensics demonstrates slightly higher watermark recovery rates than the rest semi-fragile watermarks when tested against face swapping algorithms, there remains a clear gap in the recovery rates regarding robustness and fragility. Additionally, since our watermark matrices are embedded following an entry-to-patch strategy for localizing Deepfake manipulations, only the edited patches are expected to lose the embedded watermarks. Therefore, the critically lower watermark recovery rates for face reenactment than those for face swapping are reasonable and consistent with the localization results in Section 4.4.
@@ 4.3. Deepfake Detection
Considering the robustness and fragility of watermarks regarding real and Deepfake images, semi-fragile watermarks act as a natural proactive Deepfake detector. In particular, an ideal case is expected to showcase watermark recovery rates of Deepfake images all lower than those of real images, demonstrating a clear threshold between real and fake samples. In this study, following this scheme, we evaluated the Deepfake detection performance of semi-fragile watermarking frameworks. Since no clear threshold has been provided for the comparative models, we adopted the AUC scores for evaluation [^4] The 256 resolution versions of SepMark and FractalForensics are adopted for fairness. . In addition, we included four state-of-the-art passive Deepfake detectors (Rössler et al., 2019; Shiohara and Yamasaki, 2022; Cao et al., 2022; Dong et al., 2023) that have shown reasonable performance on lab-controlled datasets for comparison. In this experiment, the real samples are derived by applying common image operations as listed in Table 2 and the fake ones are produced by adopting the Deepfake manipulations on the real images in CelebA-HQ.
@@
Figure 3. Effect of image operations on the watermarked image with localization demonstration by red overlay.
@@
As Table 4 lists, although demonstrating some acceptable detection results (e.g., SBIs (Shiohara and Yamasaki, 2022) and CADDM (Dong et al., 2023) on SimSwap), the passive detectors generally exhibit poor generalizability toward most Deepfake manipulations in the experiments, leading to average AUC scores below 70%. On the other hand, while most semi-fragile algorithms have exhibited desired watermark fragility against most Deepfake manipulations, the proactive Deepfake detection performance depends on their robustness against common image operations. As a result, since a clear gap between watermark recovery rates for real and fake can be observed regarding FractalForensics, we promisingly achieved satisfactory Deepfake detection results with an average of 99.99% AUC score over all participating Deepfake manipulations. As for the comparative ones, due to larger watermark recovery rates regarding fragility than those for robustness, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) fail to achieve Deepfake detection with poor AUC scores. SepMark (Wu et al., 2023), although shows satisfactory performance with AUC scores of 99.99% for some Deepfake models, the surprisingly high watermark recovery rates in Table 3 result in much lower AUC scores for InfoSwap (Gao et al., 2021) and E4S (Liu et al., 2023), leading to an average of 92.23% in Table 4. Meanwhile, although the watermark recovery rates are close for robustness and fragility, there exists a small gap that benefits FaceSigns (Neekhara et al., 2024) in deriving reasonable AUC scores against all Deepfake manipulations. However, its fragility against Gaussian Noise remains a potential risk.
@@
Figure 4. Effect of Deepfake manipulations on the watermarked image with localization demonstration by red overlay.
@@ 4.4. Deepfake Manipulation Localization
Localization Against Deepfake Manipulations. We then evaluated the localization results against Deepfake manipulations. Since Deepfake images are produced upon feature fusion via deep neural networks, there is no ground-truth for localizing Deepfake manipulations. Therefore, FractalForensics is designed to embed watermarks with an entry-to-patch strategy that implicitly enables the localization functionality. As a result, as visualized in Figure 4, the first and second rows display the clean Deepfake results via different manipulation algorithms and their corresponding localization results regarding the sample pair of images, respectively. In the third row, we recorded the number of 4-bit watermark entries that were lost for each patch position on the entire testing dataset and plotted the cumulative localization results after normalization. Given the fact that CelebA-HQ contains facial images that are well-aligned with faces mostly located in the center of images, it can be concluded that, since face swapping manipulations mainly modify detailed facial organs and cues, the localized manipulated areas are mainly gathered in the center of the images. It is also noticed that the localized areas for DiffSwap (Zhao et al., 2023b) are mainly in the bottom half of the face. This is because DiffSwap does not heavily edit the forehead and eyes but the nose and mouth. On the other hand, face reenactment algorithms edit not only facial expressions that are located within the face area but also the head poses that can heavily affect the background area, leading to localized areas that are largely spanned on the images. Additionally, by taking a look at Table 3 together with Figure 4, consistencies can be observed between localization visualizations and watermark recovery rates.
Furthermore, as WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) are also designed to have localization functionality against Deepfake and AIGC, we conducted comparative experiments for the models. The localization results are shown in Figure 5. Specifically, WaterLo defines purple overlays as fake areas and uses green for real, while EditGuard highlights the fake and real areas with white and black colors, respectively. However, although successfully highlighting the entire image as fake for all Deepfake-manipulated images, they both lack detailed information that highlights specific manipulated areas. On the other hand, since they each are vulnerable to some common image operations, the localization results become relatively unconvincing. Contrarily, Figures 34, and 5 prove the reliability of FractalForensics in localizing Deepfake manipulated areas and provide explainable evidence to support Deepfake detection.
Table 5. Cross-dataset evaluation on LFW for watermark recovery rate against common and visual quality.
@@
Watermark Sensitivity Against Cropping. In this study, since the watermarking pipeline follows an entry-to-patch strategy, as a side contribution, our approach also owns sensitivity against malicious cropping. To explore the watermark robustness in this case, we conducted experiments to randomly crop adjacent image patches and verify whether watermarks in the cropped areas were lost and those in the untampered areas remained. As demonstrated in Figure 6, we attempted cropping sizes from $1\times 1$ to $5\times 5$ patches, and we recorded the bit-wise and patch-wise watermark recovery rates, along with the correctness regarding desired robustness and fragility. As a result, the bit-wise and patch-wise watermark recovery rates exhibit droppings that generally match with the number of cropped patches for both CelebA-HQ and LFW, while the overall correctness stays above 95% even when 25 out of 64 image patches are cropped within an image.
@@ 4.5. Cross-Dataset Evaluation on LFW
In this section, to evaluate the cross-dataset performance of our proposed FractalForensics, we conducted experiments on LFW (Huang et al., 2012). Specifically, experiments follow the same settings as in Table 2 and Table 3 to compute the watermark recovery rates against common image operations and Deepfake manipulations. For simplicity, the 256 resolution versions for SepMark (Wu et al., 2023) and our FractalForensics are adopted in the comparison.
As Table 5 reports, the comparative watermarking frameworks generally maintain consistent performance as for CelebA-HQ (Karras et al., 2018), demonstrating obvious vulnerability when facing Gaussian Noise, with no watermark recovery rate above 90%. On the contrary, although still slightly affected by the Resize operation, our proposed FractalForensics generally preserves robustness against all common image operations at an average of 99.08% with good visual quality. Additionally, while showing outstanding visual quality, WaterLo fails to achieve the expected watermark robustness. Besides the unsatisfactory robustness and a promising PSNR value, EditGuard exhibits an unexpectedly low SSIM value, which may be caused by mismatched image color or tone.

### v2

@@ 3.4. Loss Functions
Table 2. Watermark recovery rate against common image operations on CelebA-HQ.
@@
We train the decoder for watermark robustness following
@@ 4.1. Implementation Details
We used the popular facial image datasets, CelebA-HQ (Karras et al., 2018) and LFW (Huang et al., 2012), for experiments. Specifically, we followed the official split of CelebA-HQ with 30,000 facial images for training, validation, and testing, and we used LFW with 5,749 facial identities for evaluating the generalizability. We trained two versions of FractalForensics for the 256 and 512 resolutions, with $N_{\textrm{img}}$ being 5 and 6, $N_{\textrm{wtm}}$ being 5 and 4, $N_{\textrm{rec}}$ being 5 and 4, and $N_{\textrm{dec}}$ being 4 and 3, respectively. For the loss weights, we set $\lambda_{\textrm{MSE}}=1$, $\lambda_{\textrm{LPIPS}}=0.5$, $\lambda_{\textrm{adv}}=0.01$, and $\lambda_{\textrm{dec}}=12$. During training, we used Identity, Jpeg, and Gaussian Noise as common image processing operations to boost watermark robustness. Deepfake manipulations are completely omitted during training under the black-box setting. In the testing phase, the semi-fragile watermarking algorithms are evaluated against common image processing operations including Identity, Jpeg, Gaussian Noise, Gaussian Blur, Median Blur, and Resize and Deepfake manipulations including SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), UniFace (Xu et al., 2022), E4S (Liu et al., 2023), and DiffSwap (Zhao et al., 2023b) for face swapping and StarGAN (Choi et al., 2020), StyleMask (Bounareli et al., 2023b), and HyperReenact (Bounareli et al., 2023a) for face reenactment. We trained the main watermarking framework at a learning rate of $0.002$ and trained the discriminator at $0.0004$. All experiments are conducted on 8 Tesla A100 GPUs.
@@ 4.2. Watermarking Performance on CelebA-HQ
In this section, we evaluate the watermarking performance on CelebA-HQ (Karras et al., 2018), with respect to the visual quality of the watermarked images and the watermark robustness and fragility with respect to common image operations and Deepfake manipulations.
Visual Quality. We conducted quantitative evaluations of the visual quality after embedding watermarks. Specifically, we compared our proposed FractalForensics to state-of-the-art semi-fragile watermarks for proactive defense against Deepfake and AIGC. As Table 1 reports, our method achieves the second-best performance with the most difficult bit-per-pixel rate, while the best WaterLo (Beuve et al., 2023) has watermarks with only 1 bit. In general, while all methods demonstrate reasonable LPIPS values, SepMark [^3] SepMark consists of a robust and a semi-fragile watermark, where we adopted the semi-fragile watermark for comparing robustness and fragility in this study. (Wu et al., 2023) and FaceSigns (Neekhara et al., 2024) both have PSNR values below 35, and EditGuard (Zhang et al., 2024) derives an unsatisfactory SSIM value below 0.85. As a result, our approach consistently maintains promising visual qualities at both resolutions for all evaluation metrics on visual quality.
Table 3. Watermark recovery rate against Deepfake image manipulations on CelebA-HQ.
@@
Watermark Recovery Rate. We evaluated the watermark recovery rate of our approach against common image processing operations and Deepfake manipulations, compared to contrastive semi-fragile watermarking frameworks. As demonstrated in Table 2, while all algorithms stay robust as expected when applied with no operation (Identity), they each suffer a drastic reduction in watermark recovery rates when facing some image processing operations. In general, Gaussian Noise has the most serious effect, causing all watermark recovery rates to fall below 65% for the comparative algorithms. Meanwhile, the two models that have demonstrated relatively better visual quality in Table 1, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024), also fail to maintain robustness regarding Jpeg, Gaussian Blur, Median Blur, and Resize operations. For our proposed FractalForensics, since a 4-bit binary value represents each watermark entry, we computed the patch-wise watermark recovery rate as our official performance and used that for Deepfake detection and localization in later sections, which is harder than the bit-wise metric in preserving robustness. As a result, our method achieves the best performance on average with 99.55% and 99.73% at the 256 and 512 resolutions. In the last two columns, we also presented the bit-wise watermark recovery rates of FractalForensics as references, where reasonably higher values than the patch-wise ones can be observed as expected since the patch-wise watermark recovery rate is computed following a more strict rule.
@@
When tested against Deepfake manipulations, the watermark fragility is expected to be observed. In other words, there is expected to be an obvious gap between the watermark recovery rates for robustness and fragility, such that real and fake samples are distinguishable accordingly. For this experiment, we adopted five face swapping (Chen et al., 2020; Gao et al., 2021; Xu et al., 2022; Liu et al., 2023; Zhao et al., 2023b) and three face reenactment (Choi et al., 2020; Bounareli et al., 2023b, a) manipulations for evaluation. It can be observed from Table 3 that the fragility has been achieved by WaterLo (Beuve et al., 2023), FaceSigns (Neekhara et al., 2024), EditGuard (Zhang et al., 2024), and our proposed FractalForensics, while SepMark (Wu et al., 2023) demonstrates robust watermark recovery ability from images generated by SimSwap (Chen et al., 2020), InfoSwap (Gao et al., 2021), and E4S (Liu et al., 2023). Meanwhile, FractalForensics is observed to demonstrate promising fragility at both resolutions with average watermark recovery rates of 50.05% and 39.27%.
@@
With a closer look at Table 2 and Table 3, since there exist watermark recovery rates against common image operations that are lower than those against Deepfake manipulations for the comparative models (e.g., Sepmark has 99.94% for InfoSwap but 54.47% for Gaussian Noise), it becomes impossible to distinguish real and fake based on robustness and fragility. As a result, although our FractalForensics demonstrates slightly higher watermark recovery rates than the other semi-fragile watermarks when tested against face swapping algorithms, there remains a clear gap in the recovery rates regarding robustness and fragility. Additionally, since our watermark matrices are embedded following an entry-to-patch strategy for localizing Deepfake manipulations, only the edited patches are expected to lose the embedded watermarks. Therefore, the critically lower watermark recovery rates for face reenactment than those for face swapping are reasonable and consistent with the localization results in Section 4.4.
@@ 4.3. Deepfake Detection
Considering the robustness and fragility of watermarks regarding real and Deepfake images, semi-fragile watermarks act as a natural proactive Deepfake detector. In particular, an ideal case is expected to showcase watermark recovery rates of Deepfake images all lower than those of real images, demonstrating a clear threshold between real and fake samples. In this study, following this scheme, we evaluated the Deepfake detection performance of semi-fragile watermarking frameworks. Since no clear threshold has been provided for the comparative models, we adopted the AUC scores for evaluation. In addition, we included four state-of-the-art passive Deepfake detectors (Rössler et al., 2019; Shiohara and Yamasaki, 2022; Cao et al., 2022; Dong et al., 2023) that have shown reasonable performance on lab-controlled datasets for comparison. In this experiment, the real samples are derived by applying common image operations as listed in Table 2 and the fake ones are produced by performing Deepfake manipulations on the real images in CelebA-HQ.
@@
Figure 3. Effect of image operations on the watermarked image with localization demonstrated by the red overlay.
@@
As Table 4 lists, although demonstrating some acceptable detection results (e.g., SBIs (Shiohara and Yamasaki, 2022) and CADDM (Dong et al., 2023) on SimSwap), the passive detectors generally exhibit poor generalizability toward most Deepfake manipulations in the experiments, leading to average AUC scores below 70%. On the other hand, while most semi-fragile algorithms have exhibited desired watermark fragility against most Deepfake manipulations, the proactive Deepfake detection performance depends on their robustness against common image operations. As a result, since a clear gap between watermark recovery rates for real and fake can be observed for FractalForensics, we promisingly achieved satisfactory Deepfake detection results with an average of 99.99% AUC score over all participating Deepfake manipulations. As for the comparative ones, due to larger watermark recovery rates regarding fragility than those for robustness, WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) fail to achieve Deepfake detection with poor AUC scores. SepMark (Wu et al., 2023), although shows satisfactory performance with AUC scores of 99.99% for some Deepfake models, the surprisingly high watermark recovery rates in Table 3 result in much lower AUC scores for InfoSwap (Gao et al., 2021) and E4S (Liu et al., 2023), leading to an average of 92.23% in Table 4. Meanwhile, although the watermark recovery rates are close for robustness and fragility, there exists a small gap that benefits FaceSigns (Neekhara et al., 2024) in deriving reasonable AUC scores against all Deepfake manipulations. However, its fragility against Gaussian Noise remains a potential risk.
@@
Figure 4. Effect of Deepfake algorithms on the watermarked image with localization demonstrated by the red overlay.
@@ 4.4. Deepfake Manipulation Localization
Localization Against Deepfake Manipulations. We then evaluated the localization results against Deepfake manipulations. Since Deepfake images are produced upon feature fusion via deep neural networks, there is no ground-truth for localizing Deepfake manipulations. Therefore, FractalForensics is designed to embed watermarks with an entry-to-patch strategy that implicitly enables the localization functionality. As a result, as visualized in Figure 4, the first and second rows display the clean Deepfake results via different manipulation algorithms and their corresponding localization results regarding the sample pair of images, respectively. In the third row, we recorded the number of 4-bit watermark entries that were lost for each patch position on the entire testing dataset and plotted the cumulative localization results after normalization. Given the fact that CelebA-HQ contains facial images that are well-aligned with faces mostly located in the center of images, it can be concluded that, since face swapping manipulations mostly modify detailed facial organs and cues, the localized manipulated areas are mainly in the center of the images. It is also noticed that the localized areas for DiffSwap (Zhao et al., 2023b) are mainly in the bottom half of the face. This is because DiffSwap does not heavily edit the forehead and eyes but the nose and mouth. On the other hand, face reenactment algorithms edit not only facial expressions that are located within the face area but also the head poses that can heavily affect the background area, leading to localized areas that largely span entire images. Additionally, by taking a look at Table 3 together with Figure 4, consistencies can be observed between localization visualizations and watermark recovery rates.
@@
Table 5. Cross-dataset evaluation on LFW for visual quality and watermark recovery rate against common image operations.
@@
Furthermore, as WaterLo (Beuve et al., 2023) and EditGuard (Zhang et al., 2024) are also designed to have localization functionality against Deepfake and AIGC, we conducted comparative experiments for the models. The localization results are shown in Figure 5. Specifically, WaterLo defines purple overlays as fake areas and uses green for real, while EditGuard highlights the fake and real areas with white and black colors, respectively. However, although successfully highlighting the entire image as fake for all Deepfake-manipulated images, they both lack detailed information that highlights specific manipulated areas. On the other hand, since they each are vulnerable to some common image operations, the localization results become relatively unconvincing. In contrast, Figures 34, and 5 show the reliability of FractalForensics in localizing Deepfake manipulated areas and provide explainable evidence to support Deepfake detection.
Watermark Sensitivity Against Cropping. In this study, since the watermarking pipeline follows an entry-to-patch strategy, as a side contribution, our approach also exhibits sensitivity against malicious cropping. To explore the watermark robustness in this case, we conducted experiments to randomly crop adjacent image patches and verify whether watermarks in the cropped areas were lost and those in the untampered areas remained. As demonstrated in Figure 6, we attempted cropping sizes from $1\times 1$ to $5\times 5$ patches, and we recorded the bit-wise and patch-wise watermark recovery rates, along with the correctness regarding desired robustness and fragility. As a result, the bit-wise and patch-wise watermark recovery rates exhibit declines that generally match with the number of cropped patches for both CelebA-HQ and LFW, while the overall correctness stays above 95% even when 25 out of 64 image patches are cropped within an image.
@@ 4.5. Cross-Dataset Evaluation on LFW
In this section, to evaluate the cross-dataset performance of our proposed FractalForensics, we conducted experiments on LFW (Huang et al., 2012). Specifically, experiments follow the same settings as in Table 2 and Table 3 to compute the watermark recovery rates against common image operations and Deepfake manipulations. For simplicity, the 256 resolution versions for SepMark (Wu et al., 2023) and our FractalForensics are utilized in the comparison.
As Table 5 reports, the comparative watermarking frameworks generally maintain consistent performance as for CelebA-HQ (Karras et al., 2018), demonstrating obvious vulnerability when facing Gaussian Noise, with no watermark recovery rate above 90%. In contrast, although still slightly affected by the Resize operation, our proposed FractalForensics generally preserves robustness against all common image operations at an average of 99.08% with good visual quality. Additionally, while showing outstanding visual quality, WaterLo fails to achieve the expected watermark robustness. Besides the unsatisfactory robustness and a promising PSNR value, EditGuard exhibits an unexpectedly low SSIM value, which may be caused by mismatched image color or tone.